# Multilevel lattice codes from Hurwitz quaternion integers

### v2

@@ Abstract
This work presents an extension of the Construction $\pi_{A}$ lattices proposed in [1], to Hurwitz quaternion integers. This construction is provided by using an isomorphism from a version of the Chinese remainder theorem applied to maximal orders in contrast to natural orders in prior works. Exploiting this map, we analyze the performance of the resulting multilevel lattice codes, highlight via computer simulations their notably reduced computational complexity provided by the multistage decoding. Moreover it is shown that this construction effectively attain the Poltyrev-limit.
@@ I Introduction
Lattices have been utilized in communication scenarios for several decades, presenting considerable benefits. Some coding techniques employed across various applications rely on a randomized collection of nested lattice codes derived from Construction A, which can achieve the capacity of the additive white Gaussian noise (AWGN) channel [2, 3]. However, it’s important to note that decoding lattices constructed via Construction A requires decoding the underlying linear codes implemented over a finite ring, leading to substantial complexity in decoding.
@@
To address this challenge, Huang and Narayanan proposed a lattice construction known as Construction $\pi_{A}$ firstly considering integers $\mathbb{Z}\subset\mathbb{R}$, Gaussian integers $\mathbb{Z}[i]\subset\mathbb{C}$, and Eisenstein integers $\mathbb{Z}[\omega]\subset\mathbb{C}$ [1]. This construction can be viewed as a special case of Construction A, applied to codes represented as the Cartesian product of $k$ linear codes. These codes are considered over $R/p_{1}R,...,R/p_{k}R$, where $R=\mathbb{Z},\mathbb{Z}[i]$ or $\mathbb{Z}[\omega]$ and $p_{i}$ is a prime number in the related ring. Such constructions are made possible through a ring isomorphism between the product of prime fields and the quotient rings $\left(R/\prod_{j=1}^{k}p_{j}R\right)$, established by the Chinese Remainder Theorem (CRT). Additionally, due to its multilevel nature, Construction $\pi_{A}$ lattices offer the advantage of multistage decoding, where each class representative is decoded level by level. Another extensions to this construction was proposed in [4] using imaginary quadratic fields and in [5] for natural orders over cyclic division algebras.
@@
Several studies have explored the utility of cyclic division algebras employing natural orders [6, 7, 8]. Utilizing maximal orders has emerged as an effective strategy, enabling denser constructions without compromising crucial parameters like minimum distance. The initial work on maximal orders to design space-time (ST) codes was proposed in [9], subsequently leading to the derivation of the densest ST codes [10]. Furthermore, various studies have explored the numerous advantages of employing maximal orders in quaternionic settings [11, 12]. Additionally, these orders offer significant advantages, particularly in multiple-input/single-output (MISO) [9] and multiple-input/multiple-output (MIMO) transmissions [13, 10].
In this paper, we introduce a novel extension of Construction $\pi_{A}$ lattices [1] utilizing the maximal order of Hurwitz quaternion integers and a modified version of the Chinese Remainder Theorem. This approach enables the generation of codes with distinct sizes from those obtained in [5]. We employ Loeliger’s approach [14] instead of Forney’s [15] as in [1] to demonstrate that the multilevel lattice codes associated with this construction attain the Poltyrev-limit for the AWGN channel, and their reduced computational complexity is also analysed through computer simulations.
@@
Recent applications in communications have exploited the mathematical structure of quaternions. In this context, the maximal order of quaternionic integers in Hurwitz’s $\mathscr{H}$ holds significance in Hamilton’s restriction of quaternion algebras, offering favorable properties for constructing residue class rings with a finite number of elements. An important relationship for such applications is the fact that $\mathscr{H}$ is a quaternion-valued representation of the $D_{4}^{*}$ lattice, providing the best lattice sphere packing efficiency in this dimension [16, 17]. For instance, the Hurwitz lattice has the double of the packing density of the lattice $\mathbb{Z}^{4}$ which is associated to the natural order of the Lipschitz quaternion integers. This property has been especially exploited in optical communications [18, 19, 20, 21]. Due to its favorable algebraic properties, particularly the possibility to define quotient rings in $\mathscr{H}$, Hurwitz quaternion integers have also been explored for broadcast transmission over the AWGN channel [22, 5] and MIMO transmissions [23, 24, 25, 26]. Notably, with a more theoretical approach, [27] demonstrates that Hurwitz lattices in dimension $4n$ also linearly improve the Minkowski-Hlawka lower bound.
@@
The remaining of this paper is organized as follows: Section II provides preliminaries on lattices codes and quaternion algebras to be used in the development of this work. Section III presents the extension of Construction $\pi_{A}$ to the Hurwitz integers. Section IV describes a decoding algorithm for Construction $\pi_{A}$ lattice over Hurwitz quaternion integers and discuss its decoding complexity. In Section V is shown that this extension can achieve the AWGN Poltyrev-limit for channel coding and in Section VI conclusions and future perspectives are drawn.
@@ II-A Lattices and lattice codes
A lattice $\Lambda$ can be defined as a discrete subgroup of $\mathbb{R}^{n}$. Equivalently, a lattice is a subset of $\mathbb{R}^{n}$ generated by all linearly integer combination of $m$ linearly independent vectors, $v_{1},...,v_{m}$. A matrix $B$ whose columns are the vectors $v_{i}$ is called a generator matrix of $\Lambda$. We deal here only with full-rank lattice ($m=n$).
@@
    For $\mathbf{x},\mathbf{y}\in\Lambda$, where $\mathbf{x}\neq\mathbf{y}$, $(\mathbf{x}+\mathscr{R})$ and $(\mathbf{y}+\mathscr{R})$ intersect at most on their boundaries and,
@@
    $\bigcup_{\mathbf{x}\in\Lambda}(\mathbf{x}+\mathscr{R})=\mathbb{R}^{n}$, i.e., the entire $\mathbb{R}^{n}$ space is tiled through translations by points of $\Lambda$.
An example of fundamental region is the fundamental parallelotope, $\mathscr{P}(B)$ defined as,
$\mathscr{P}(B)=\left\{\sum_{i=1}^{n},\alpha_{i}v_{i},0\leq\alpha_{i}<1\right\}.$
The volume of a lattice, $\text{vol}(\Lambda)=|\det(B)|$, is the volume of any fundamental region.
Another fundamental region associated to a lattice is the Voronoi region, $\mathscr{V}_{\Lambda}(\mathbf{x})$, at a point $\mathbf{x}\in\Lambda$, is the set
$\mathscr{V}_{\Lambda}(\mathbf{x})=\{\mathbf{y}\in\mathbb{R}^{n};|\mathbf{x}-\mathbf{y}|\leq|z-\mathbf{y}|,\ \forall z\in\Lambda\}.$ (1)
Given a point $z\in\mathbb{R}^{n}$ and a lattice $\Lambda\subset\mathbb{R}^{n}$, we define $Q_{\Lambda}(z)$ as the closest lattice point to $z$.
$Q_{\Lambda}(z)=\mathbf{x}\in\Lambda;\ \ ||z-\mathbf{x}||\leq||z-\mathbf{y}||\ \ \ \ \forall\ \mathbf{y}\in\Lambda.$ (2)
and ties must be chosen. Subtracting the closest lattice point $Q_{\Lambda}(z)$ from $z$ wraps the real vector $z$ into the Voronoi region at the origin, $\mathscr{V}_{\Lambda}(0)=\mathscr{V}_{\Lambda}$. This operation, called the modulo-$\mathbf{\Lambda}$ operation, is denoted as
@@
The minimum distance between two distinct points of a lattice corresponds to the minimum norm of non-zero vectors in $\Lambda$, i.e.,
$d_{\text{min}}=\displaystyle\min_{0\neq\mathbf{x}\in\Lambda}\|\mathbf{x}\|.$ (4)
The sphere packing of a lattice $\Lambda$ as the union of all the translated balls with radius $\rho=d_{\text{min}}/2$ centred at points of $\Lambda$, $\rho$ is called the packing radius of $\Lambda$.
A linear code over $\mathbb{Z}/q\mathbb{Z}=\mathbb{Z}_{q}$, the ring of integer modulo $q$, is a subset $\mathscr{C}\subset\mathbb{Z}_{q}^{n}$ closed under addition.
A method for constructing lattices from linear codes is the well known Construction A. A Construction A lattice is usely defined from linear codes over $\mathbb{Z}_{q}$ and also from codes over $\mathbb{Z}[i]$ and $\mathbb{Z}[\omega]$, number fields or division algebras [16, 29, 30].
@@
    Consider the linear code $\mathscr{C}=\{\mathbf{x}=G\cdot\mathbf{y}:\mathbf{y}\in\mathbb{Z}_{q}^{k}\}$, where all operations are over $\mathbb{Z}_{q}$.
@@
    $\Lambda_{A}(\mathscr{C})=\{\mathbf{x}\in\mathbb{Z}^{n}:\mathbf{x}\mod q\in\mathscr{C}\}=\mathscr{C}+q\mathbb{Z}^{n}.$
It is shown that $\Lambda_{A}(\mathscr{C})$ is a full-rank lattice, $q\mathbb{Z}^{n}\subset\Lambda_{A}\subset\mathbb{Z}^{n}$, and that the volume of this lattice is $q^{n}/M$, where $M$ is the size of the code $\mathscr{C}$ [16].
In [14], an important result (due Hlawka) of the Minkowski-Hlawka theorem [31, 32] was used to show that randomly selecting a code from the random $(n,k)$-linear code over $\mathbb{Z}_{p}$, $p$ prime, and applying the Construction A produces lattices that are suitable for channel coding as $p$ tends to infinity.
Consider the communication channel $\mathbf{y}=\mathbf{x}+\mathbf{w}$, where $\mathbf{y}$ is the received signal, $\mathbf{x}\in\Lambda$ is the transmitted signal with no power constraint, and each element in $\mathbf{w}$ is i.i.d. $\sim\mathscr{N}(0,\sigma^{2})$. Consider a target error probability $10<\epsilon<$ 1, the normalized volume to noise ratio (NVNR) of a lattice $\Lambda\subset\mathbb{R}^{n}$ is defined as
$\mu(\Lambda,\epsilon)=\dfrac{\text{vol}(\Lambda)^{2/n}}{\sigma^{2}(\epsilon)}.$
@@
Then, a sequence of lattices $\Lambda_{n}$ of increasing dimension is Poltyrev-good [33] if,
@@ II-B Quaternion Algebras
Let $\mathbb{F}$ be a field. An algebra $B$ over $\mathbb{F}$ (or $\mathbb{F}-$algebra) is a vector space over $\mathbb{F}$ equipped with associative multiplication
@@
$(x,y)$ $\mapsto xy$
satisfying $(k_{1}x+k_{2}y)z=k_{1}(xz)+k_{2}(yz)\ \text{and}\ x(k_{1}y+k_{2}z)=k_{1}(xy)+k_{2}(xz)$ for all $k_{1},k_{2}\in\mathbb{F}$, $x,y,z\in B$.
@@
Let $\mathbb{F}$ be a field with characteristic different from 2. A quaternion algebra over $\mathbb{F}$ is an $\mathbb{F}$-algebra admitting a basis of four elements, denoted $1,i,j,k$, which satisfy the following relations: $1$ is the neutral element for multiplication, and
@@
for some elements $a,b\in\mathbb{F}^{\times}$. We denote this algebra by $(a,b)_{\mathbb{F}}$.
The classical example to be considered here is the quaternion algebra over the reals due to Hamilton [35, (1848)],
@@
If $B$ is assumed to be a quaternion algebra over the field $\mathbb{F}=\mathbb{Q}$ of rational numbers with $a,b<0$, $B=(a,b)_{\mathbb{Q}}$. $B$ is also called a definite quaternion algebra.
@@
An order $O\subseteq B=(a,b)_{\mathbb{Q}},a,b<0$ is a full $\mathbb{Z}$-lattice that is also a subring having $1\in B$.
@@
Considering $B=(a,b)_{\mathbb{Q}}$ a definite quaternion algebra, we have
@@
is an order.
@@
    $\lambda\mathbf{x}\in I$ for $\lambda\in O$ and $\mathbf{x}\in I$.
We can define the right ideals analogously and we call a two-sided ideal any left ideal that is also a right ideal.
@@
We define $\mathfrak{p}\subsetneq O$ as left prime ideal whenever $\mathfrak{a},\mathfrak{b}\in O$ are left ideals such that $\mathfrak{a}\mathfrak{b}\in\mathfrak{p}$ implies that either $\mathfrak{a}\in\mathfrak{p}$ or $\mathfrak{b}\in\mathfrak{p}$.
@@
A two-sided ideal $\mathfrak{P}\subseteq O\subset B$ is said to be a prime ideal if $\mathfrak{P}\neq O$, and for two-sided ideals $\mathfrak{U},\mathfrak{W}\subseteq O$, we have,
[…]
@@
An order $O\subseteq B=(a,b)_{\mathbb{Q}},a,b<0$ is maximal if it is not properly contained in another order of $B$.
@@
=$\mathscr{L}=\{a_{1}1+a_{2}i+a_{3}j+a_{4}k|a_{1},...,a_{4}\in\mathbb{Z}\},$ (6)
where $(1,i,j,k)$ is the standard basis of $B$. A direct verification shows that $\mathscr{L}$ is an order of $B$ called the Lipschitz order. This is not a maximal order.
The maximal order in $B=(-1,-1)_{\mathbb{Q}}$ is given by the Hurwitz quaternion integers [38] defined next.
@@
The Hurwitz quaternion integers (Hurwitz integers), $\mathscr{H}\subset\mathbb{H}$, is defined as the set,
=$\mathscr{H}=\{a_{1}+a_{2}i+a_{3}j+a_{4}\left(\frac{1+i+j+k}{2}\right)\ |\ a_{1},...,a_{4}\in\mathbb{Z}\}.$
From now on we restrict our work to $B=(-1,-1)_{\mathbb{Q}}$ and the maximal order $\mathscr{H}$ of Hurwitz integers. Note that $\mathscr{H}$ is composed by all quaternions having each component as an integer or a “half-integer”.
@@
For an element $\alpha=a+bi+cj+dk\in\mathscr{H}$, its conjugated $\overline{\alpha}$ is $a-bi-cj-dk$. The norm of $\alpha$ is $\text{Nrm}(\alpha)=\alpha\cdot\overline{\alpha}=\overline{\alpha}\cdot\alpha=a^{2}+b^{2}+c^{2}+d^{2}\in\mathbb{Z}$ and $Re(\alpha)=a$ represents the real part of $\alpha$. An element $\varepsilon$ is a unit iff $\text{Nrm}(\varepsilon)=1$. The set of units, $\mathscr{H}^{\times}$ is composed for $24$ units, namely,
×={±$\mathscr{H}^{\times}=\{\pm 1,\pm i,\pm j,\pm k,\dfrac{\pm 1\pm i\pm j\pm k}{2}\}.$
The set of Hurwitz integers has some important properties, such as a left Euclidean algorithm which can be obtained as in the commutative case.
Using left division, we can define the left greatest common divisor and also a version of Bézout’s identity holds:
@@
Note that in $\mathbb{Z}[i]$ and $\mathbb{Z}[\omega]$ the concept of irreducible corresponds to "prime" as in the natural numbers. But in $\mathscr{H}$ we may have an irreducible quaternion $\pi$, to divide from the left a product $\alpha\beta$, $\pi\rceil\alpha\beta$, without having either $\pi\rceil\alpha$ or $\pi\rceil\beta$. For example, let $\delta=(1+j+k)(1+j)=-i+2j+k$ and $\pi=1+i-j$, then $\pi\rceil\delta$ since can write $\delta=(1+j)\pi$ but we do not have either $\pi\rceil(1+j+k)$ or $\pi\rceil(1+j)$.
For Hurwitz integers, it is shown that the irreducible elements are those with a norm equal to a rational prime.
@@
[…]
@@
A result that provides the construction of multilevel lattice codes is the Chinese remainder theorem (CRT). Inspired by the Construction $\pi_{A}$ proposed in [1, 4] and [5], we found the following version of CRT that can be applied to maximal orders over definite quaternion algebras and therefore to the Hurwitz integers.

### v3

@@ Abstract
This work presents an extension of the Construction $\pi_{A}$ lattices proposed by Huang and Narayanan, to Hurwitz quaternion integers. This construction is provided by using an isomorphism from a version of the Chinese remainder theorem applied to maximal orders, in contrast to natural orders in prior works. Exploiting this map, we analyse the performance of the resulting multilevel lattice codes and highlight via computer simulations their notably reduced computational complexity provided by the multistage decoding. Moreover, it is shown that there is a sequence of Construction $\pi_{A}$ lattices that attain with high probability the Poltyrev-limit.
Index Terms:
Lattice codes, multilevel decoding, quaternion algebras, Chinese remainder theorem.
@@ I Introduction
Lattices have been used in communication scenarios for several decades, regarding coding for reliable and secure transmission through different channels. Some coding techniques employed rely on a randomized collection of nested lattice codes derived from Construction A, which can achieve the capacity of the additive white Gaussian noise (AWGN) channel [1, 2]. However, decoding Construction A lattices typically involves decoding the underlying linear codes over a finite ring, leading to substantial computational complexity in decoding.
@@
To address this challenge, Huang and Narayanan proposed in [3] a multilevel lattice construction known as Construction $\pi_{A}$ firstly considering integers $\mathbb{Z}\subset\mathbb{R}$, Gaussian integers $\mathbb{Z}[i]\subset\mathbb{C}$, and Eisenstein integers $\mathbb{Z}[\omega]\subset\mathbb{C}$. This construction can be viewed as a special case of Construction A, applied to codes represented as the Cartesian product of $r\in\mathbb{N}$ linear codes. These codes are considered over the residue class rings $R/p_{1}R,\ldots,R/p_{k}R$, where $R=\mathbb{Z},\mathbb{Z}[i]$ or $\mathbb{Z}[\omega]$ and $p_{i}$ is a prime number in the related ring. Such constructions are made possible through a ring isomorphism between the product of finite rings and the quotient rings $\left(R/\prod_{j=1}^{r}p_{j}R\right)$, established by the Chinese Remainder Theorem (CRT). Additionally, due to its multilevel nature, Construction $\pi_{A}$ lattices offer the advantage of multistage decoding, where each class representative is decoded level by level. Further extensions of this construction were proposed in [4] using imaginary quadratic fields and in [5] for natural orders over cyclic division algebras. In [6], it is proposed a scheme using partitions of the golden code, where it can be viewed as a space-time (ST) code derived from a cyclic division algebra. The authors addressed the difficulty of applying the CRT in non-commutative settings by embedding the cyclic division algebra into a commutative structure, enabling partitions of the golden code.
@@
In the context of communication theory, cyclic division algebras, particularly through the use of natural and maximal orders, have been proved useful in the design of ST codes. Several studies have explored natural orders for explicit constructions of fully diverse and fast-decodable ST codes [7, 8, 9]. Later, the use of maximal orders has emerged as an effective strategy, enabling denser constructions without compromising crucial parameters like minimum distance. The initial work on maximal orders to design ST codes was proposed in [10], subsequently leading to the derivation of the densest ST codes [11]. Furthermore, various studies have explored the numerous advantages of employing maximal orders in quaternionic settings [12, 13]. These orders offer significant advantages, particularly in multiple-input/single-output (MISO) [10] and multiple-input/multiple-output (MIMO) transmissions [14, 11].
@@
A particularly rich structure arises from the Hurwitz quaternion integers $\mathscr{H}$, considered here, which forms a maximal order in the Hamilton quaternion algebra $\mathbb{H}$. The quotient rings of the format $\mathscr{H}/p\mathscr{H}$ (p odd prime) provide code constructions with favourable properties. An important relationship regarding applications is the fact that $\mathscr{H}$ is a quaternion-valued representation of the $D_{4}^{*}$ lattice, providing the best lattice sphere packing efficiency in this dimension [15, 16]. For instance, the Hurwitz lattice (the lattice associated to $\mathscr{H}$) has the double of the packing density of the lattice $\mathbb{Z}^{4}$ which is associated to the natural order of the Lipschitz quaternion integers. These properties have motivated its use in optical communications [17, 18, 19, 20], broadcast transmission over the AWGN channel [21, 5] and MIMO transmissions [22, 23, 24, 25]. Notably, with a more theoretical approach, it was shown that Hurwitz lattice in dimension $4n$ improve the Minkowski-Hlawka lower bound [26], highlighting their strong potential in lattice coding.
In this paper, we propose a novel extension of Construction $\pi_{A}$ lattices [3] utilizing the maximal order of Hurwitz quaternion integers and a modified version of the CRT adapted to this non-commutative structure. This approach enables the construction of multilevel codes with different sizes than those obtained in [5] and supports efficient multistage decoding. We adopt Loeliger’s averaging approach [27] instead of Forney’s [28] as in [3] to demonstrate that the resulting lattice achieves the Poltyrev-limit over the AWGN channel, and to analyse their computational complexity via numerical simulations. We also include a preliminary investigation of Hurwitz-based Construction $\pi_{A}$ lattice codes in the context of index coding problems, where their algebraic structure may offer advantages in side information exploitation.
@@
The remaining of this paper is organized as follows: Section II provides the necessary background on lattice codes and quaternion algebras. Section III presents the extension of Construction $\pi_{A}$ to Hurwitz integers. Section IV describes a decoding algorithm for Construction $\pi_{A}$ lattices over the Hurwitz quaternion integers and discusses its decoding complexity. In Section V it is shown that this extension can achieve the Poltyrev-limit for channel coding. Section VI presents an example of the use of this setting to index coding problem, and in Section VII conclusions and future perspectives are drawn.
@@ II-A Lattices and lattice codes
A lattice $\Lambda$ can be defined as an additive discrete subgroup of $\mathbb{R}^{n}$. Equivalently, a lattice is a subset of $\mathbb{R}^{n}$ generated by all integer linear combinations of $m$ independent vectors, $v_{1},\ldots,v_{m}\in\mathbb{R}^{n}$. A matrix $B$ whose columns are the vectors $v_{i}$ is called a generator matrix of $\Lambda$. We deal here only with full-rank lattices ($m=n$).
@@
    For $x,y\in\Lambda$, where $x\neq y$, $(x+\mathscr{R})$ and $(y+\mathscr{R})$ intersect at most on their boundaries and,
@@
    $\bigcup_{x\in\Lambda}(x+\mathscr{R})=\mathbb{R}^{n}$, i.e., the entire $\mathbb{R}^{n}$ space is tiled through translations by points of $\Lambda$.
An example of fundamental region of $\Lambda$ is the fundamental parallelotope, $\mathscr{P}(B)$ associated to a generator matrix $B$ and defined as,
$\mathscr{P}(B)=\left\{\sum_{i=1}^{n}\alpha_{i}v_{i},0\leq\alpha_{i}<1,\alpha_{i}\in\mathbb{R}\right\}.$
The volume of a lattice $\Lambda$, $\text{vol}(\Lambda)=|\det(B)|$, is the volume of any fundamental region of $\Lambda$.
Another fundamental region associated to a lattice $\Lambda$ is the Voronoi region, $\mathscr{V}_{\Lambda}(x)$, at a point $x\in\Lambda$:
$\mathscr{V}_{\Lambda}(x)=\{y\in\mathbb{R}^{n};|x-y|\leq|\lambda-y|,\ \forall\lambda\in\Lambda\}.$ (1)
Given a point $z\in\mathbb{R}^{n}$ and a lattice $\Lambda\subset\mathbb{R}^{n}$, we define $Q_{\Lambda}(z)$ as the closest lattice point to $z$ as
$Q_{\Lambda}(z)=\lambda\in\Lambda;\ \ ||z-\lambda||\leq||z-\hat{\lambda}||\ \ \ \ \forall\ \hat{\lambda}\in\Lambda,$ (2)
where ties are broken systematically as in [25], for example. Subtracting the closest lattice point $Q_{\Lambda}(z)$ from $z$ wraps the real vector $z$ into the Voronoi region at the origin, $\mathscr{V}_{\Lambda}(0)=\mathscr{V}_{\Lambda}$. This operation, called the modulo-$\Lambda$ operation, is denoted as
@@
The minimum distance between two distinct points of a lattice $\Lambda$ corresponds to the minimum Euclidean norm of non-zero vectors in $\Lambda$, i.e.,
$d_{\text{min}}(\Lambda)=\displaystyle\min_{0\neq x\in\Lambda}\|x\|.$ (4)
The sphere packing of a lattice $\Lambda$ is the union of all the translated balls with radius $\rho=d_{\text{min}}(\Lambda)/2$ centred at points of $\Lambda$ and $\rho$ is called the packing radius of $\Lambda$.
A linear code over $\mathbb{Z}/q\mathbb{Z}=\mathbb{Z}_{q}$, the ring of integers modulo $q$, is a subset $\mathscr{C}\subset\mathbb{Z}_{q}^{n}$ closed under addition.
A method for constructing lattices from linear codes is the well known Construction A. A Construction A lattice is usually defined from linear codes over $\mathbb{Z}_{q}$, but can also be considered from codes over $\mathbb{Z}[i]$ and $\mathbb{Z}[\omega]$, number fields or division algebras [15, 30, 31].
@@
    Consider the linear code $\mathscr{C}=\{x=G\cdot y:y\in\mathbb{Z}_{q}^{k}\}$, where all operations are over $\mathbb{Z}_{q}$.
@@
    $\Lambda_{A}(\mathscr{C})=\{x\in\mathbb{Z}^{n}:x\mod q\in\mathscr{C}\}=\mathscr{C}+q\mathbb{Z}^{n}.$
It is shown that $\Lambda_{A}(\mathscr{C})$ is a full-rank lattice, $q\mathbb{Z}^{n}\subset\Lambda_{A}(\mathscr{C})\subset\mathbb{Z}^{n}$, and that the volume of this lattice is $q^{n}/M$, where $M$ is the size of the code $\mathscr{C}$ [15].
In [27], an important result of the Minkowski-Hlawka theorem [32, 33] was used to show that by randomly selecting a code from the set of all $(n,k)$-linear code over $\mathbb{Z}_{p}$, $p$ prime, Construction A produces lattices that are suitable for channel coding as $p$ tends to infinity.
Consider the unconstrained AWGN channel $y=x+z$, where $y$ is the received signal, $x\in\Lambda$ is the transmitted signal with no power constraint, and each element in $z$ is i.i.d. $\sim\mathscr{N}(0,\sigma^{2})$. Consider a target error probability $10<\epsilon<$ 1, the normalized volume to noise ratio (NVNR) of a lattice $\Lambda\subset\mathbb{R}^{n}$ is defined as
$\mu(\Lambda,\epsilon)=\dfrac{\text{vol}(\Lambda)^{2/n}}{\sigma^{2}(\epsilon)},$
@@
Then, a sequence of lattices $\Lambda_{n}$ of increasing dimension is Poltyrev-good [34] if
@@ II-B Quaternion Algebras
Let $\mathbb{K}$ be a field. An algebra $B$ over $\mathbb{K}$ (or $\mathbb{K}-$algebra) is a vector space over $\mathbb{K}$ equipped with an associative multiplication
@@
$(x,y)$ $\mapsto xy,$
satisfying $(k_{1}x+k_{2}y)z=k_{1}(xz)+k_{2}(yz)\ \text{and}\ x(k_{1}y+k_{2}z)=k_{1}(xy)+k_{2}(xz)$ for all $k_{1},k_{2}\in\mathbb{K}$, $x,y,z\in B$.
@@
Let $\mathbb{K}$ be a field with characteristic different from 2. A quaternion algebra over $\mathbb{K}$ is an $\mathbb{K}$-algebra admitting a basis of four elements, denoted $1,i,j,k$, which satisfy the following relations: $1$ is the neutral element for multiplication, and
@@
for some non-vanishing elements $a,b\in\mathbb{K}$. We denote this algebra by $(a,b)_{\mathbb{K}}$.
The classical example to be considered here is the quaternion algebra over the reals due to Hamilton [36, (1848)],
@@
If $B$ is assumed to be a quaternion algebra over the field $\mathbb{K}=\mathbb{Q}$ of rational numbers with $a,b<0$, $B=(a,b)_{\mathbb{Q}}$, $B$ is also called a definite quaternion algebra.
@@
An order $O\subset B=(a,b)_{\mathbb{Q}},a,b<0$ is a full $\mathbb{Z}$-lattice that is also a subring having $1\in B$.
@@
Consider $B=(a,b)_{\mathbb{Q}}$ a definite quaternion algebra. Then
@@
is an order in $B$, called a natural order.
Due to the non-commutativity of multiplication, it is necessary to distinguish between left ideals, right ideals and two-sided ideals. Throughout this paper, we focus only on left operations, although equivalent results can be obtained for right operations as well.
@@
    $\lambda x\in I$ for any $\lambda\in O$ and $x\in I$.
Right ideals are defined analogously, and a two-sided ideal is a subset that is both a left and a right ideal.
@@
We define $\mathfrak{p}\subsetneq O$ a left prime ideal whenever $\mathfrak{a},\mathfrak{b}\subset O$ are left ideals such that $\mathfrak{a}\mathfrak{b}\subset\mathfrak{p}$ implies that either $\mathfrak{a}\subset\mathfrak{p}$ or $\mathfrak{b}\subset\mathfrak{p}$.
@@
A two-sided ideal $\mathfrak{P}\subsetneq O\subset B$ is said to be a prime ideal if, for two-sided ideals $\mathfrak{U},\mathfrak{W}\subset O$, we have,
$\mathfrak{U}\cdot\mathfrak{W}\subset\mathfrak{P}\Rightarrow\mathfrak{U}\subset\mathfrak{P}\ \ \textrm{or}\ \ \mathfrak{W}\subset\mathfrak{P}.$
@@
An order $O\subset B=(a,b)_{\mathbb{Q}},a,b<0$ is maximal if it is not properly contained in another order of $B$.
@@
=$\mathscr{L}=\{a_{1}1+a_{2}i+a_{3}j+a_{4}k|a_{1},\ldots,a_{4}\in\mathbb{Z}\},$ (6)
where $(1,i,j,k)$ is the standard basis of $B$. A direct verification shows that $\mathscr{L}$ is an order of $B$ called the Lipschitz order. This is not a maximal order [39].
The maximal order in $B=(-1,-1)_{\mathbb{Q}}$ is given by the Hurwitz quaternion integers [39], defined next.
@@
The Hurwitz quaternion integers (Hurwitz integers), $\mathscr{H}\subset\mathbb{H}$, is defined as the set
=$\mathscr{H}=\left\{a_{1}+a_{2}i+a_{3}j+a_{4}\left(\frac{1+i+j+k}{2}\right)\ |\ a_{1},\ldots,a_{4}\in\mathbb{Z}\right\}.$
From now on, we restrict our work to $B=(-1,-1)_{\mathbb{Q}}$ and the maximal order $\mathscr{H}$ of the Hurwitz integers. Note that $\mathscr{H}$ is composed by all quaternions having each component as an integer or a “half-integer”.
@@
For an element $\alpha=a+bi+cj+dk\in\mathscr{H}$, its conjugated $\overline{\alpha}$ is $a-bi-cj-dk$. The norm of $\alpha$ is $\text{Nrm}(\alpha)=\alpha\cdot\overline{\alpha}=\overline{\alpha}\cdot\alpha=a^{2}+b^{2}+c^{2}+d^{2}\in\mathbb{Z}$ and $Re(\alpha)=a$ represents the real part of $\alpha$. An element $\varepsilon$ is a unit iff $\text{Nrm}(\varepsilon)=1$. The set of units, $\mathscr{H}^{\times}$, is composed of $24$ units, namely
×={±$\mathscr{H}^{\times}=\left\{\pm 1,\pm i,\pm j,\pm k,\dfrac{\pm 1\pm i\pm j\pm k}{2}\right\}.$
The set of Hurwitz integers has some important properties, such as admitting a left Euclidean algorithm which can be obtained as in the commutative case [39].
Using left division, we can define the left greatest common divisor, and also a version of Bézout’s identity holds.
@@
Note that in $\mathbb{Z}[i]$ and $\mathbb{Z}[\omega]$ the concept of irreducible corresponds to “prime” as in the natural numbers. But in $\mathscr{H}$ we may have an irreducible quaternion $\pi$ to divide from the left a product $\alpha\beta$, $\pi\rceil\alpha\beta$, without having either $\pi\rceil\alpha$ or $\pi\rceil\beta$. For example, let $\delta=(1+j+k)(1+j)=-i+2j+k$ and $\pi=1+i-j$, then $\pi\rceil\delta$ since we can write $\delta=(1+j)\pi$ but we do not have either $\pi\rceil(1+j+k)$ or $\pi\rceil(1+j)$.
For Hurwitz integers, it is shown that the irreducible elements are those with norm equal to a rational prime.
@@
Let $\pi\in\mathscr{H}$. Then $\pi$ is irreducible if and only if $\textrm{Nrm}(\pi)=p\in\mathbb{Z}$ is a rational prime.
@@
A result that provides the construction of multilevel lattice codes is the CRT. Inspired by the Construction $\pi_{A}$ proposed in [3, 4] and [5], we found the following version of CRT that can be applied to maximal orders over definite quaternion algebras and therefore to the Hurwitz integers.